OpenStudy (anonymous):

solving for x 8800= 70x10^9 * 0.00025 + x/3.25 + 200x10^9 * 0.00013 * x/ 5.30 I know it's suppose to be a small number, something to the -4 power.

OpenStudy (anonymous):

Mathematica 8 was used for the solution of x. 8800= 70x10^9 * 0.00025 + x/3.25 + 200x10^9 * 0.00013 * x/ 5.30 was pasted into a notebook as "Free-form input" and became 8800 = 70000000000*0.00025 + x/3.25 + 200000000000*0.00013*(x/5.3) Following an automatic simplification, the equation was changed to: \[8800=1.75\times 10^7+4.90566\times 10^6 x \] Solving the expression above for x gave the following result to 18 digits of precision: x = -3.5655136225180803 Verifing the computed value for x: The 8000 on the left side of the original equation was moved to the right hand side and evaluated at x = -3.5655136225180803. The result was " {0.} " to an estimated 16 digits of precision.

OpenStudy (anonymous):

8000 in the last paragraph should have been typed in as 8800.